OpenStudy (anonymous):

Solve the problem below and write out your steps. Then, use your work to write a detailed explanation of the steps you took to solve the problem. Make sure you use proper terminology. 3(7t + -3) = -4t + 41

OpenStudy (alexwee123):

well you would first distribute out the 3 so 21t-9=-4t+41

OpenStudy (alexwee123):

then add 4t and 9 on both sides 25t=50 divide by 25 t=2

OpenStudy (anonymous):

\[21t-9=-4t+41\]\[21t+4t=41+9\]\[25t=50\]\[t=2\]